Write an equation and solve.
Josef has a jar of dimes,nickels, and pennies. He has one dime more than five times the number of nickels and four times as many pennies as nickels. How many of each coin does he have if their total value is $7.77?

Answer by macston(5194)

.
N=nickels; P=pennies=4N; D=dimes=5N+1
.
1(P)+5(N)+10(D)=777
1(4N)+5N+10(5N+1)=777
4N+5N+50N+10=777
59N=767
N=13 ANSWER 1: He has 13 nickels.
P=4N=4(13)=52 ANSWER 2: He has 52 pennies.
D=5N+1=5(13)+1=66 ANSWER 3: He has 66 dimes
.
CHECK:
$0.01P+$0.05N+$0.10D=$7.77
$0.01(52)+$0.05(13)+$0.10(66)=$7.77
$0.52+$0.65+$6.60=$7.77
$7.77=$7.77